Shift Summary for owl shift, 27/aug/99

AG (ag@fnal.gov)
Fri, 27 Aug 1999 08:05:57 -0400 (EDT)

Shift Summary for owl shift, 27/aug/99

On Shift: BobT, AG

Hours of beam available at all intensities: 4h
Hours of beam available at nominal intensity: 3h
Hours of E832 data (special runs not included): 2.5h
Hours of special runs (laser scans, muon runs...): 4h

Accounting of lost time:

No beam. Confusion with new tapes loaded.

========================================================================

Runs ended this shift:

Run 14340 - COSMIC
Run 14344 - MUON, DC plateau run.

Tapes brought to FCC this shift:
(including tapes for online split)

None.

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

No beam till after 3am. When beam returns we make muon run to make DC
plateaus, and then start taking e832 data.

========================================================================

Description of Shift:

run____spill______time_______comment

------------------------------------------------------------------------

14340____________0030______Taku walks in and scientific discussion
___________________________about the future of pi0nunubar Dalitz
___________________________mode fills the control room.

_________________0100______Taku makes a move to leave, but
___________________________then decides to stay and brings up a
___________________________theological paladrome dilemma:
___________________________"God sees dog." and "Dog sees God."...
___________________________in which religions are both statements valid?
_________________0113______Shift crew shares its ideas with Taku and he
___________________________leaves quite satisfied, but with a strong
___________________________desire to leave for Japan very early
___________________________in the morning...

_________________0309______MCR calls and promises low intensity beam soon.

_________________0312______Stop COSMIC run.

_________________0317______Beam! 2.5e12 . We page Hogan, who comes in
___________________________shortly.

_________________0320______Initializing MUON run. All tapes come up red
___________________________and we decide to make dart fresh start.
___________________________This doesn't help. We page Sydney, who comes
___________________________in to discover that these are new tapes!!
___________________________That was not clear from logbook or run log or
___________________________run summary.

_________________0355______We start run 14344 which is a plateau of
___________________________chambers run w/ different user latch settings
___________________________for different voltages, details in logbook.

14345____________0518______We start BEAM_832 run and beam promptly goes
___________________________away.

_________________0540______Beam returns.

========================================================================

Follow up Notes on Previous Shift

o Hogan paged and DC plateaus are done during the muon run.
o Tapes are labeled

========================================================================

Items to Follow up on Next Shift